### v1

# London, UK Dual Branch VideoMamba for Violence Detection

Abstract
The rapid proliferation of surveillance cameras has increased the demand for automated violence detection. While CNNs and Transformers have shown success in extracting spatio-temporal features, they struggle with long-term dependencies and computational efficiency. We propose Dual Branch VideoMamba with Gated Class Token Fusion (GCTF), an efficient architecture combining a dual-branch design and a state-space model (SSM) backbone where one branch captures spatial features, while the other focuses on temporal dynamics, with continuous fusion via a gating mechanism. We also present a new benchmark by merging RWF-2000, RLVS, and VioPeru datasets in video violence detection, ensuring strict separation between training and testing sets. Our model achieves state-of-the-art performance on this benchmark offering an optimal balance between accuracy and computational efficiency, demonstrating the promise of SSMs for scalable, real-time surveillance violence detection.
1 Introduction
Figure 1: (a) Accuracy vs FLOPS, (b) Accuracy vs Parameters.
The increased installation of surveillance cameras in public and private spaces driven by advancements in low-cost imaging technology has led to a dramatic increase in the amount of surveillance video data generated daily [Ullah et al.(2023)Ullah, Obaidat, Ullah, Muhammad, Hijji, and Baik, Omarov et al.(2022)Omarov, Narynov, Zhumanov, Gumar, and Khassanova]. Since continuous human monitoring of these numerous surveillance video feeds is neither practical nor reliable to detect violent behaviors, there is a significant interest in developing deep learning models that can reliably and efficiently detect violence in surveillance videos.
Recent advancements in computer vision have leveraged deep convolutional neural networks (CNNs) and, more recently, Transformer-based architectures to extract rich spatio-temporal features from video data [Ji et al.(2012)Ji, Xu, Yang, and Yu, Arnab et al.(2021)Arnab, Dehghani, Heigold, Sun, Lučić, and Schmid]. However, while 3D CNNs excel at capturing local patterns, they often struggle with the long-term dependencies that are crucial for understanding extended video sequences. On the other hand, Transformer models, although powerful in modeling global context, suffer from quadratic computational complexity, which limits their scalability [Arnab et al.(2021)Arnab, Dehghani, Heigold, Sun, Lučić, and Schmid].
In this work, we propose the Dual Branch VideoMamba with Gated Class Token Fusion (GCTF) model which is designed to address these limitations with the help of an efficient state-space model (SSM) backbone. Based on the recent VideoMamba framework [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao], we introduce a dual-stream processing architecture that continuously fuses the learned information from both branches through a gated class token fusion mechanism between the two branches. Our model processes two parallel streams: one dedicated to capturing fine-grained spatial cues and the other focused on temporal information while fusing learned features of each branch through fusing the CLS tokens pertaining to each branch continuously with the help of a gated learnable parameter. This fusion allows the network to effectively detect violent activities even in scenarios where subtle motion differences are critical.
In addition to the novel architectural design, we address a key challenge in surveillance violence detection dataset diversity and generalizability by introducing an amalgamated dataset that combines three established surveillance benchmarks: Real World Fighting (RWF-2000) [Cheng et al.(2021)Cheng, Cai, and Li], Real Life Violence Situations (RLVS) [Soliman et al.(2019)Soliman, Kamal, Nashed, Mostafa, Chawky, and Khattab], and VioPeru [Huillcen Baca et al.(2024)Huillcen Baca, Palomino Valdivia, and Gutierrez Caceres]. This integrated surveillance dataset provides a broader range of violent scenarios, improving the generalizability of our approach across different environments and recording conditions. Our contributions are threefold:
  1. 1.
    We propose a dual-branch VideoMamba architecture that leverages the efficiency of state-space models while explicitly decoupling spatial and temporal reasoning through two specialized pipelines with different scanning strategies. Unlike prior dual-encoder SSM models, our architecture is tailored for classification tasks without memory modules, and enables cross-branch interaction at the semantic (class token) level.
  2. 2.
    We introduce a novel continuous Gated Class Token Fusion (GCTF) mechanism that enables adaptive integration of the class tokens from the dual streams, ensuring a balanced and adaptive fusion of learned spatial and temporal features.
  3. 3.
    We curate an integrated surveillance violence detection dataset by amalgamating RWF-2000, RLVS, and VioPeru, while rigorously preventing data leakage between training and testing sets where extensive experiments demonstrate that our Dual Branch VideoMamba with GCTF model achieves state-of-the-art performance in surveillance violence detection with respect to the combined dataset while increasing computational efficiency as evident in Fig. 1 and Table 2, making it a promising solution for real-world surveillance applications.
2 Related Work
Research in automated violence detection has primarily followed two paradigms: anomaly detection and action recognition. In the anomaly detection paradigm [Pu et al.(2024)Pu, Wu, Yang, and Wang, Chen et al.(2023)Chen, Liu, Zhang, Fok, Qi, and Wu], violent events are modeled as rare deviations from normal activity. While, this approach has shown promise in controlled environments, they have often failed to capture the complex context in which violence occurs in real-world [Sultani et al.(2018)Sultani, Chen, and Shah, Chen et al.(2023)Chen, Liu, Zhang, Fok, Qi, and Wu, Tian et al.(2021)Tian, Pang, Chen, Singh, Verjans, and Carneiro].
In contrast, the action recognition paradigm treats violence detection as a supervised classification problem where the focus is on learning discriminative features for violent versus non-violent actions. Early methods utilized 3D CNNs to capture spatio-temporal features directly from video clips [Ullah et al.(2019)Ullah, Ullah, Muhammad, Haq, and Baik]. Subsequent research [Su et al.(2020)Su, Lin, Zhu, and Wu] introduced techniques such as skeleton-based action recognition, where human pose estimation is used to generate 3D skeleton representations, and Graph CNNs are employed to model the interactions between individuals. [Su et al.(2020)Su, Lin, Zhu, and Wu] is one of the first papers to evaluate performance on a real-world surveillance violence detection data set (RWF-2000) [Cheng et al.(2021)Cheng, Cai, and Li] where almost all the previous literature were evaluated on non-surveillance based datasets such as the Hockey Fight dataset [Bermejo Nievas et al.(2011)Bermejo Nievas, Deniz Suarez, Bueno García, and Sukthankar]. [Garcia-Cobo and SanMiguel(2023), Asad et al.(2022)Asad, Jiang, Yang, Tu, and Malik, Islam et al.(2021)Islam, Rukonuzzaman, Ahmed, Kabir, and Farazi] have used deep architectures comprising of two simultaneous pipelines to extract different types of features similarly to [Feichtenhofer et al.(2019)Feichtenhofer, Fan, Malik, and He] in action recognition space. These methods have shown robust performance on benchmark datasets, but still face challenges in effectively modeling long-term dependencies due to their reliance on standard convolutional operations.
Lately, Transformer-based models have been applied to violence detection by leveraging self-attention mechanisms to capture global context [Li et al.(2023a)Li, Yang, and Liang]. Recently, models such as [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh] have incorporated architectures that amalgamate convolution and self-attention mechanisms to balance computational efficiency with representational power to effectively capture local and global dependencies in the context of violence detection inspired by video action recognition models such as [Li et al.(2023b)Li, Wang, He, Li, Wang, Wang, and Qiao]. Although such models have achieved impressive performance, their quadratic complexity in attention operations can hamper their scalability [Li et al.(2023c)Li, Wang, Zhang, Gao, Song, Liu, Li, and Qiao, Li et al.(2023b)Li, Wang, He, Li, Wang, Wang, and Qiao].
The emergence of efficient subquadratic-time architectures with state-space models (SSMs) has opened a new avenue for addressing these challenges [Gu and Dao(2024), Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. VideoMamba [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao] extends the work of [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] in the video recognition space where it replaces the traditional self-attention mechanism with a linear-complexity state-space module, enabling efficient processing of video sequences without sacrificing the ability to capture long-range dependencies. Similarly, [Li et al.(2024b)Li, Zhao, Yang, Liu, Sheng, Zeng, Wang, Wu, and Jiang] has introduced another SSM architecture for video anomaly detection by effectively modeling spatial-temporal normality through a Spatial-Temporal interaction module. Our approach mitigates the drawbacks of traditional CNNs and Transformers while benefiting from the computational efficiency of SSMs. To our knowledge, this paper is the first to investigate the use of SSMs for video violence detection.
3 Proposed Method
Our design builds upon the VideoMamba framework and its efficient state-space model modules. We employ two parallel pipelines—each implementing a variant of VideoMamba with distinct scanning strategies—with the intention of separately extracting spatial and temporal features from video inputs, combined with a cropping module to focus on human-interactions. We introduce a novel Gated Class Token Fusion (GCTF) mechanism that combines information between the two branches. GCTF is performed at each layer in the network, providing a form of continuous fusion. The enriched feature representation is combined from the Spatial-First Scanning Branch to the Temporal-First Scanning Branch to form a unified representation for final classification.
3.1 Overview of Mamba Architectures
The Mamba family of architectures rethinks the standard self-attention mechanism by leveraging efficient state-space models (SSMs) for long-range dependency modeling  [Gu and Dao(2024)]. The key idea is to view a sequence as the output of a continuous dynamical system, where the evolution of a hidden state is governed by ordinary differential equations [Gu and Dao(2024)]. In its continuous form, a state-space model is given by:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t),$ (1)
where $x(t)$ is the input at time $t$, $h(t)\in\mathbb{R}^{N}$ is the hidden state, and where $\mathbf{A}\in\mathbb{R}^{N\times N}$ represents the evolutionary matrix of the system and $\mathbf{B}\in\mathbb{R}^{N\times D}$, and $\mathbf{C}\in\mathbb{R}^{d\times N}$ are projection matrices. To process discrete token sequences, this continuous system is approximated via discretization (commonly using a zero-order hold), which includes a timescale parameter $𝚫{\mathbf{\Updelta}}$ to transform the continuous parameters ${\mathbf{A}},{\mathbf{B}}$ to their discrete learnable counterparts $\overline{{\mathbf{A}}},\overline{{\mathbf{B}}}$:
[…]
$h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},\quad y_{t}=\mathbf{C}h_{t}.$ (3)
[…]
3.2 Dual Branch VideoMamba with GCTF Architecture
We introduce our novel architecture, the Dual Branch VideoMamba with Gated Class Token Fusion (GCTF) for Violence Detection in surveillance videos as shown in Fig. 2. The architecture contains four main components, namely: (a) Cropping Module; (b) Branch-1 (Spatial-First Scanning); (c) Branch-2 (Temporal-First Scanning); and (d) Final Fusion Block; inspired by the motivational factors discussed in the preceding section.
Figure 2: Part (a): The overall Dual Branch VideoMamba with GCTF Architecture with: 1) Cropping module to detect people and crop spatially the area of focus; 2) Branch-1 and Branch-2 which implement the VideoMamba architecture in parallel with lateral connections for continuous Gated Class Token Fusion (GCTF); 3) Final Fusion Block. Part (b): Bi-Directional Vision Mamba Encoder Block [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang].
3.2.1 Cropping Module
To focus on human actions, our cropping mechanism extracts the region that encompasses all detected people in each frame to guide the network, based on the observation that violent incidents typically involve multiple individuals. More details on this are discussed in Section 6.1 of supplementary material. This approach, also adopted in CUE-Net [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh], enhances the model’s ability to accurately identify violent behavior as evident in the ablation study.
3.2.2 Dual-Branch Architecture Overview
[…]
$\mathbf{X}^{1}={}$ $\left[\mathbf{X}_{cls}^{1},\mathbf{X}^{1}\right]+\mathbf{p}_{s}^{1}+\mathbf{p}_{t}^{1},$ (4)
$\mathbf{X}^{2}={}$ $\left[\mathbf{X}_{cls}^{2},\mathbf{X}^{2}\right]+\mathbf{p}_{s}^{2}+\mathbf{p}_{t}^{2},$ (5)
and subsequently are fed into the two distinct pipelines.
Figure 3: (a) Spatial-First Scanning (b) Temporal-First Scanning.
Branch-1 (Spatial-First Scanning Pipeline):
This branch reorganizes the tokens such that the spatial layout within each frame is prioritized. In other words, tokens are grouped and ordered by their spatial coordinates before being concatenated across frames as depicted in Fig. 3 (a). This strategy is aimed at learning fine-grained local spatial features that are critical for recognizing visual cues. This Spatial-First scan in this branch is performed bi-directionally.
Branch-2 (Temporal-First Scanning Pipeline):
Here, the tokens are reorganized based on the frame and then stacked along the spatial dimension maintaining their natural temporal order so that the sequential progression of frames is preserved as depicted in Fig. 3 (b). This ordering can facilitate the model to capture dynamic motion patterns and global temporal dependencies pertaining to each of those tokens, which are essential for detecting violent actions and changes over time. Similar to the Branch-1, this scan too is performed bi-directionally.
This strategy enables leveraging the pre-trained weights of a vanilla VideoMamba to initialize each branch in our architecture. Pre-trained initialization is beneficial because it provides a robust starting point, accelerates convergence during training, and enhances overall performance by transferring learned representations from extensive prior training  [Li et al.(2023c)Li, Wang, Zhang, Gao, Song, Liu, Li, and Qiao]. During the forward pass, Branch-1 processes its input video to produce a sequence of features, including a dedicated class (CLS1) token for each layer. Meanwhile, Branch-2 receives two inputs: the video embedding tokens padded with the CLS2 token and additional intermediate (CLS1) token features from each block of Branch-1 (excluding the final block). This second input allows Branch-2 to incorporate spatial context learned from Branch-1 into its feature learning to produce better results as evident by the results in Table 5 of the ablation study.
3.2.3 Gated Class Token Fusion (GCTF) between Branch-1 and Branch-2
After processing each block in the network, both pipelines output their respective CLS tokens:
  • $\text{CLS}_{l}^{1}\in\mathbb{R}^{d}$ from the spatial-first scanning Branch-1,
  • $\text{CLS}_{l}^{2}\in\mathbb{R}^{d}$ from the temporal-first scanning Branch-2.
To combine these complementary features, a learnable gate is applied between each parallel block. The gating mechanism is defined as:
$\sigma^{\prime}_{l}={}Sigmoid(\sigma_{l})$ (6)
$\text{CLS}_{\text{fused(l)}}^{2}=\sigma^{\prime}_{l}\odot\text{CLS}_{l}^{2}+\left(1-\sigma^{\prime}_{l}\right)\odot\text{CLS}_{l}^{1},$ (7)
where $\sigma_{l}\in\mathbb{R}^{d}$ is a learnable parameter vector for block $l$, and it is passed through the Sigmoid function to ensure that the gate value lie between 0 and 1 before performing element wise multiplication denoted by $\odot$ with CLS tokens from each branch to dynamically weigh the best contributions. Unlike typical fusion schemes that operate at a single depth or combine all tokens, GCTF selectively merges class-level semantics across branches at every depth, enabling layerwise refinement and preventing early overcommitment to either modality.
3.3 Final Fusion Block
At the very end of the architecture, a final fusion block integrates the final CLS tokens from Branch-1 and Branch-2 through concatenation. The output $\mathbf{Z}\in\mathbb{R}^{2d}$ is obtained as:
$\mathbf{Z}=\operatorname{concat}(CLS_{fin}^{1},CLS_{fin}^{2}),$ (8)
Finally, the target class Pr is obtained by passing Z through a fully connected projection layer.
4 Experiments and Results
4.1 Datasets
The most challenging datasets so far in the real-world violence detection domain are the Real-World Fighting (RWF-2000) dataset [Cheng et al.(2021)Cheng, Cai, and Li], Real Life Violence Situations (RLVS) dataset [Soliman et al.(2019)Soliman, Kamal, Nashed, Mostafa, Chawky, and Khattab], and VioPeru dataset [Huillcen Baca et al.(2024)Huillcen Baca, Palomino Valdivia, and Gutierrez Caceres] that contain video footage of fighting in real life scenarios. Out of these three datasets, while the RWF-2000 and VioPeru datasets contain exclusive surveillance footage, RLVS contains violence action which includes surveillance footage as well. All three datasets contain videos cropped at 5s and depict real-world scenarios. We considered other violence-related datasets, however, we observed these have many clips which are not strongly labeled or acquired from staged violence. Therefore, for the rest of the paper, we focus on these three data sets with the intention of assessing our model to reflect real-world violence scenarios.
4.1.1 Combined Dataset of RWF-2000, RLVS, and VioPeru
In order to create a more comprehensive and challenging benchmark for violence detection, we combine the three datasets: RWF-2000, RLVS, and VioPeru. Although each dataset has its own predefined train-test split, simply merging them without ensuring strict partition integrity can lead to potential data leakage, where identical videos appear in both the training and testing sets, since RWF-2000 and RLVS were both sourced from YouTube  [Cheng et al.(2021)Cheng, Cai, and Li, Soliman et al.(2019)Soliman, Kamal, Nashed, Mostafa, Chawky, and Khattab]. To address this concern, we embed all videos from the training and testing splits of each dataset using VideoMAE  [Tong et al.(2022)Tong, Song, Wang, and Wang] embeddings. Next, we compute the cosine similarity between every pair of videos (across training and testing splits) to detect duplicates. All the pairs which yielded 78% or higher cosine similarity were manually inspected.
Final Combined Dataset Statistics: Duplicates in the training sets were not investigated because they naturally represent the frequency of common real-world situations, ensuring the model learns an accurate distribution. Also, since regularization techniques are used in our model training, we can safely assume that biased learning from duplicate data is mitigated. After eliminating the identified duplicates from the RLVS test split, the statistics of final combined dataset is reported in Table 1.
Training Set Testing Set
Violent Non-Violent Violent Non-Violent
RWF-2000 800 800 200 200
RLVS 800 800 200 200
VioPeru 112 112 28 28
(Duplicates) * * (-1) 0
Total 1712 1712 427 428
Table 1: Summary of the combined dataset after removing duplicate entries.
By merging these three datasets while enforcing a strict data leakage check, we obtain a unified benchmark that provides diverse scenarios of violent and non-violent events, captured in varying resolutions and environmental conditions. This amalgamated dataset thus serves as a more robust testbed for evaluating the generalizability and performance of violence detection models.
Implementation Details: Our algorithm was implemented with the specifications of the VideoMamba-Middle architecture [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao] with 32 layers where the hidden dimension $d$ was 576 across the layers, for each branch in order to facilitate loading pre-trained weights of the vanilla VideoMamba trained with Kinetics-400 dataset [Kay et al.(2017)Kay, Carreira, Simonyan, Zhang, Hillier, Vijayanarasimhan, Viola, Green, Back, Natsev, et al.]. Further details can be found in the Section 6.4 of supplementary material.
4.2 Results
Arch. Model Pretraining Input Size Params (M) FLOPS (G) Top-1 Acc. (%) F1-Viol.(%) F1-Non-Viol.(%)
CNN SlowFast - 64$\times$2242 60 234 76.37 74.56 77.95
Trans. VideoSwin-B K-400 64$\times$2242 88 281.6 80.47 79.15 81.63
CNN + Trans. Uniformer-V2 K-400 64$\times$2242 354 6108 92.51 92.87 92.12
SSM VideoMamba K-400 64$\times$2242 74 806 94.39 94.29 94.48
CNN + Trans. CUE-Net K-400 64$\times$2242 354 5826 95.91 95.95 95.86
SSM Dual Branch VideoMamba K-400 64$\times$2242 154.3 1612 96.37 96.33 96.42
Table 2: Comparative results among various architectures for the newly combined dataset. Each model is characterized by its backbone type, the pretraining dataset, the input resolution, the number of parameters, the FLOPS, the Top-1 accuracy (%), and the F1-scores (%) for violent and non-violent classes.
In this section we perform an in-depth analysis comparing our Dual Branch VideoMamba architecture with other leading architectures using the combined dataset. Following the standard practice used for the three combined datasets, we also use classification accuracy as the main metric to evaluate the performance. The results in Table 2 show a clear progression in performance across the listed architectures. While the SlowFast model [Feichtenhofer et al.(2019)Feichtenhofer, Fan, Malik, and He], VideoSwin Transformer [Liu et al.(2022)Liu, Ning, Cao, Wei, Zhang, Lin, and Hu] and UniFormer-V2 [Li et al.(2023b)Li, Wang, He, Li, Wang, Wang, and Qiao] demonstrate respectable accuracies, they are outperformed by VideoMamba [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao] and CUE-Net [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh]. Notably, our Dual Branch VideoMamba model attains the highest scores, with a test accuracy of 96.37% and corresponding F1-scores of 96.33% and 96.42% for violence and non-violence, respectively. Furthermore, these top accuracies are achieved with a model size of 154 M parameters and a FLOPS count of 1612 GFLOPS, compared to the CUE-Net model which has 354 M parameters and 5826 GFLOPS. This represents more than a 50% reduction in both the number of parameters and the FLOPS count. Such improvements highlight the efficiency of our dual-branch SSM design over traditional CNN+Transformer architectures. Performance of our architecure on each individual dataset is detailed in Section 6.3 of supplementary material.
Figure 4: Visual Analysis of Class Activation Maps (CAMs). (a) CAMs for two violent-labelled videos correctly classified as violent. (b) CAMs for two non-violent videos misclassified as violent focusing on a forceful object removal and a collision.
Visual Analysis of Results: We use Grad-CAM [Selvaraju et al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and Batra] to visualize class activation maps (CAMs). As SSM models motion through internal state evolution rather than spatial saliency, temporal interpretability remains a key challenge. So we analyze CAMs from the spatial branch to partially address this. The spatial-first branch effectively highlights object-level cues via spatial gradients in Mamba blocks. In Fig. 4(a), the CAMs for two violent-labeled videos that were correctly classified as violent, show that the model accurately focuses on regions where human interactions which are indicative of violence occur, even when these interactions are partially occluded. In contrast, Fig. 4(b) shows CAMs for two non-violent videos that the model mistakenly classified as violent due to a forceful object removal and an accidental collision. These observations indicate that, while the model has effectively learned to detect violent cues, it may occasionally overemphasize certain visual signals too.
4.3 Ablation Study
We performed a series of ablation studies to assess the efficacy of the components of our architecture. Table 4 compares model performance with and without the cropping module. We observe that the architecture benefit from the cropping mechanism, as it focuses the network on the regions where people are most likely to appear. In Table 4, we investigate the optimal configuration for lateral connections between the two branches. Continuous lateral connections (i.e., applying them at every block) yield the highest accuracy of 96.37% where this finding highlights the importance of gradually fusing information from both branches continuously throughout the network rather than restricting fusion to early or late stages. In Table 5, we compare 3 fusion mechanisms for lateral connections: gated, additive, and concatenation-based approaches where Gated Lateral Connections (from Branch-1 to Branch-2) outperform all other methods. These results indicate that the Gated Class Token Fusion (GCTF) mechanism is highly effective in adaptively blending complementary features between branches and also shows that the direction of information flow from spatial-first scanning to temporal-first scanning also plays a critical role in maximizing performance.
Model Cropping Module Accuracy (%)
VideoMamba $\times$ 94.39
VideoMamba $\checkmark$ 95.79
Dual Branch VideoMamba $\times$ 95.44
Dual Branch VideoMamba $\checkmark$ 96.37
Table 3: Performance comparison with and without the cropping module.
Lateral-Connection-(LC) Freq. Accuracy (%)
LCs alternatively (even layers) 95.78
LCs alternatively (odd layers) 95.78
One LC only at beginning 95.44
One LC only at end 95.67
One LC at middle 95.56
Two LCs at the end and beginning 95.91
Continuous LCs 96.37
Table 4: Performance for different configurations of lateral connections.
Fusion Mechanism Accuracy (%)
Additive LCs 94.78
Concatenated LCs 95.91
Gated LCs (Branch-2 to Branch-1) 94.04
Gated LCs (Branch-1 to Branch-2) 96.37
Table 5: Performance comparison for different fusion mechanisms.
5 Conclusion
In this paper, we present a novel Dual Branch VideoMamba architecture with Gated Class Token Fusion (GCTF) for violence detection in surveillance based videos. Our method integrates an efficient state-space model with a dual-stream design aimed at separately capturing spatial features and temporal dynamics. By continuously fusing the class tokens from both branches using a learnable gating mechanism, our approach effectively combines complementary cues, yielding state-of-the-art performance. Extensive experiments on the combined benchmark dataset demonstrate that our model not only achieves state-of-the-art accuracies but also shows a significant reductions in model parameters and computational cost in FLOPS compared to previous state-of-the-art methods such as CUE-Net [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh]. This work highlights the promise of state-space models for scalable and real-time video violence detection in surveillance and paves the way for future research, including the integration of multi-modal data and further refinement of the fusion strategies.
Acknowledgment. This work has been funded through an EPSRC DTP studentship at Queen Mary University of London in collaboration with Remark AI UK Ltd. This paper utilised Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT Services [King et al.(2017)King, Butcher, and Zalewski].